# Collapse of the superconducting order parameter in Ising superconductors with Rashba spin-orbit coupling

### v1

@@ Abstract
Introduction.— Ising superconductors (SC) are doped two-dimensional transition metal dichalcogenides (TMDs), which are characterized by strong resilience against in-plane magnetic fields beyond the Pauli-paramagnetic limit. As a platform, TMDs are under active interest due to their direct band gap atomic scale thickness and strong spin-orbit coupling (SOC) [manzeli_2d_2017]. Like graphene, TMDs have a honeycomb lattice and hence two valleys. Unlike graphene, however, TMDs exhibit strong SOC because of their heavy atoms and in-plane inversion symmetry breaking. These features make them particularly interesting for potential applications in optoelectronics, valleytronics, and spintronics [xiao_coupled_2012, mak_control_2012]. When sufficiently doped, TMDs can become superconducting (SC) [kuzmanovic_tunneling_2022, de_la_barrera_tuning_2018, hamill_two-fold_2021, simon_transition-metal-dichalcogenide_2024, ilic_enhancement_2017, mockli_ising_2020, ramires_tailoring_2018, mockli_ising_2020, mockli_magnetic-field_2019, costanzo_tunnelling_2018]. Due to their large SOC, Cooper pairs are Ising coupled, making them resistant to large external magnetic fields [kuzmanovic_tunneling_2022, de_la_barrera_tuning_2018, hamill_two-fold_2021, simon_transition-metal-dichalcogenide_2024, ilic_enhancement_2017, mockli_ising_2020, ramires_tailoring_2018, mockli_ising_2020, mockli_magnetic-field_2019, costanzo_tunnelling_2018].
@@
Besides the Ising SOC, one expects a Rashba SOC to originate from gating or interfacial effects in experiment—due to the heavy atoms of the Ising SC. The effect of Rashba SOC can lead to new physics. For instance, previously, it has been shown that the presence of Rashba SOC can lead to topological superconducting phases [yuan_possible_2014, meidan_josephson_junction_2024]. Experimentally, the strength of Rashba SOC has been claimed to be of the order of the SC order parameter [lu_evidence_2015, lu_full_2018], making Rashba SOC possibly important for understanding the phase diagram of Ising SCs.
@@
In this Letter, we show that Rashba SOC sets a critical magnetic field—in the $s$-wave two valley model of Ising SCs—beyond which the superconducting order parameter (SC OP) collapses. This effect occurs at low temperatures and disappears close to the critical temperature $T_{c}$. While the effect of Rashba SOC is usually subtle, e.g., in the superconducting diode effect [yuan_topological_2021, ilic_theory_2022, khodas_diode_effect_Ising_2025]. This breakdown of the superconducting gap is a rather strong effect. Our results show that low and high temperature physic of SCs can behave both quantitatively and qualitatively different. Furthermore, our results may lead to new ways to determine the strength of the Rashba SOC in Ising SCs, which is currently a challenging task.
@@
The model.— The low energy physics of doped transition metal dichalcogenides is described by two valleys around the $\mathrm{K}$ and $\mathrm{K}^{\prime}=-\mathrm{K}$ points and a gamma pocket $\Gamma$. In the proceeding, we focus on the physics of the $\pm\mathrm{K}$ points and ignore the physics of the Gamma pocket. The general Hamiltonian near the $\pm\mathrm{K}$ [xiao_coupled_2012, kormanyos_k_2015, mockli_ising_2020, ilic_enhancement_2017] points in the $(a_{\mathbf{p},\uparrow},a_{\mathbf{p},\downarrow})$ basis reads
@@
The off-diagonal term $\mathrm{i}\sigma_{y}\otimes\delta\check{H}$ in the inverse Green’s function above, can be treated perturbatively since $|\Delta|/\sqrt{H^{2}+\beta^{2}}$ is found to be at most of the order $\mathcal{O}(10^{-1})$ in Ising SCs [de_la_barrera_tuning_2018, falson_type-ii_2020, hamill_two-fold_2021, xi_ising_2016, kuzmanovic_tunneling_2022, lu_full_2018, lu_evidence_2015]. Hence, to good accuracy we may expand the Gorkov Green’s function up to first order. The lowest order Gorkov Green’s function has a BCS form and reads
@@
$\frac{f_{\mathrm{i}\omega_{n},+,-}^{0}(\eta,\phi)}{\bar{H}\Delta}=$ $\sum_{\nu}\frac{2\pi\mathrm{i}\mathcal{N}\mathrm{sign}(\omega_{n})/2}{\mathrm{i}\omega_{n}-\sqrt{H^{2}+\beta^{2}}},$
$\frac{f_{\mathrm{i}\omega_{n},-,+}^{0}(\eta,\phi)}{-\bar{H}\Delta}=$ $\sum_{\nu}\frac{2\pi\mathrm{i}\mathcal{N}\mathrm{sign}(\omega_{n})/2}{\mathrm{i}\omega_{n}+\sqrt{H^{2}+\beta^{2}}}.$
@@
In the following, we discuss the effect of Rashba SOC on the magnetic field dependence of the SC OP at low temperatures. In principle, the effect of Rashba on Ising SCs is two-fold. One effect is that Rashba SOC coupling alters the thermodynamic ground state to a helical state $\Delta\rightarrow\Delta_{0}e^{\mathrm{i}qx}$ [kaur_helical_2005, yuan_topological_2021]. This effect is however negligible with respect to the magnitude of the SC OP since $v_{F}q$ is at most of the order $\mathcal{O}(\alpha p_{F}/\beta)$ and the correction of the helical phase to $|\Delta|/\Delta_{0}$ is at most of the order $\mathcal{O}(\alpha^{2}p_{F}^{2}/\beta^{2})<\mathcal{O}(10^{-2})$. However, the main effect of Rashba SOC to the magnitude of the SC OP is that the SC OP collapses beyond a critical magnetic field set by Rashba SOC, as we will show next.
@@
In the following, we focus on low temperatures $k_{B}T\ll\alpha p_{F}\lesssim\Delta_{0}\ll\sqrt{H^{2}+\beta^{2}}$. Translated to the critical temperature, this means $T/T_{c}\lesssim 10^{-2}$. Since $T_{c}\sim\mathcal{O}(10\mathrm{K})$ in Ising SCs, this corresponds to temperatures of the order $T\lesssim\mathcal{O}(100\mathrm{mK})$ [kuzmanovic_tunneling_2022, hamill_two-fold_2021, li_recent_2021, xi_ising_2016]. Such temperatures are typically accessible in experiment [kuzmanovic_tunneling_2022]. Furthermore, we consider the energy scale of Rashba SOC to be at most of the order of the energy scale of the SC OP at zero temperature. As detailed in the supplementary material, we perform the summation over Matsubara frequencies in the first term of Eq. 3, which is known to give the digamma function [kopnin_theory_2009] and approximate the second term for low temperatures. As a result, the self-consistency equation (3) for low temperatures reads
@@
In Refs. [lu_evidence_2015, lu_full_2018] values of $\alpha p_{F}\sim\Delta_{0}$ in $\mathrm{MoS_{2}}$ and $\mathrm{WS_{2}}$ were reported. This should be sufficient for the collapse of the SC OP to happen at low temperatures.
@@
Hence, the effective gap that is measured via the density of states is $\bar{\beta}|\Delta|$ rather than $|\Delta|$ and Rashba SOC makes the density of states cone-shaped for energies $\epsilon$ satisfying $|\epsilon-\bar{\beta}|\Delta||<\alpha p_{F}H$, see Fig. 4. Moreover, inclusion of the second order correction in Gorkov Green’s function (2) gives what is called a mirage gap at high energies $\epsilon\sim\sqrt{\beta^{2}+H^{2}}$ in the density of states [tang_magnetic_2021].
@@
Discussion.— In this Letter, we showed that the presence of Rashba SOC in Ising superconductors can strongly affect the magnetic field dependence of the superconducting order parameter at low temperatures. Our main finding is that the SC OP collapses beyond a critical magnetic field satisfying $\alpha p_{F}H=\beta|\Delta|$. This collapse disappears at high temperatures—close to $T_{c}$—and demonstrates how the low- and high temperature limits of Ising SCs can be quantitatively and qualitatively different. We find Rashba SOC to be significant to the superconducting phase diagram if $\alpha p_{F}/\Delta_{0}\gtrsim\mathcal{O}(10^{-1})$. Due to the heavy atom nature of TMDs, such a Rashba SOC could be induced from gating or interfacial effects in experiments. Experimentally, Rashba SOCs of magnitude of the SC OP have already been measured in thin-film Ising SCs [lu_evidence_2015, lu_full_2018]. Our findings may lead to new ways to determine the size of Rashba SOC in Ising SCs, which is currently a challenging task.
The collapse of the SC OP at low temperatures should be experimentally verifiable using tunneling spectroscopy [kuzmanovic_tunneling_2022, dvir_spectroscopy_2018, costanzo_tunnelling_2018]. While different tunneling devices have different sensitivities towards either the $\Gamma$ pocket or the $\pm\mathrm{K}$ valleys, the predicted effect should be within experimental reach using tunneling devices that are mostly sensitive to the density of states in the $\pm\mathrm{K}$ valleys [kuzmanovic_tunneling_2022].
@@
In future work, one could extend this model to include the $\Gamma$ pocket. Additionally, inclusion of impurities and scattering between the $\pm\mathrm{K}$ valleys and $\Gamma$ pocket would be of interest.
@@ Acknowledgements
I Self-consistency equation in the low temperature limit
The self consistency equation from the original basis to the rotated basis is given by
$\frac{\Delta}{\lambda}=$ $\frac{k_{B}T}{2}\sum_{n}\int\frac{\mathrm{d}\phi}{2\pi}\mathrm{Tr}\left\{f_{\mathrm{i}\omega_{n}}\mathrm{i}\sigma_{y}\right\}\rightarrow\frac{k_{B}T}{2}\sum_{n}\int\frac{\mathrm{d}\phi}{2\pi}\mathrm{Tr}\left\{f_{\mathrm{i}\omega_{n}}\left(\mathcal{U}_{\mathbf{-k}}^{T}\mathrm{i}\sigma_{y}\mathcal{U}_{\mathbf{k}}\right)\right\}.$ (6)
Filling in the anomalous Green’s function, we find the self-consistency equation (3), which reads
$\frac{1}{\mathcal{N}\lambda}=\pi\mathrm{i}k_{B}T\sum_{n,\nu}$ $\int\frac{\mathrm{d}\phi}{2\pi}\left[\frac{\bar{\beta}^{2}}{\mathrm{i}\sqrt{\left(\omega_{n}+\mathrm{i}\big[\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F,\nu}\big]\right)^{2}+\bar{\beta}^{2}|\Delta|^{2}}}+\frac{\bar{H}^{2}\mathrm{sign}(\omega_{n})}{\mathrm{i}\omega_{n}-\nu\sqrt{H^{2}+\beta^{2}}}\right].$ (7)
We furthermore assume $k_{B}T\ll\alpha p_{F}\lesssim\Delta_{0}\ll\sqrt{H^{2}+\beta^{2}}$ and notice that the sum over Matsubara frequencies of the left term can be transformed into an integral in real space
$\frac{1}{\mathcal{N}\lambda}=$ (8)
$\sum_{\nu}\int\frac{\mathrm{d}\epsilon}{4}\int\frac{\mathrm{d}\phi}{2\pi}\;\frac{\bar{\beta}^{2}\tanh\left[{\left(\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F,\nu}\right)}/{2k_{B}T}\right]}{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}+\pi\mathrm{i}k_{B}T\sum_{n,\nu}\Bigg[\frac{\bar{H}^{2}\mathrm{sign}(\omega_{n})}{\mathrm{i}\omega_{n}-\nu\sqrt{H^{2}+\beta^{2}}}\Bigg].$
In order to perform the integral in Eq. 8, we use the identity $\tanh(x)=1-2n_{F}(x)$. By noticing that $1/\mathcal{N}\lambda=\bar{\beta}^{2}/\mathcal{N}\lambda+\bar{H}^{2}/\mathcal{N}\lambda$, we may write the first line in Eq. 8 as
$\frac{1}{\mathcal{N}\lambda}+\sum_{\nu}\int\frac{\mathrm{d}\epsilon}{4}\frac{1}{\sqrt{\epsilon^{2}+\bar{\beta}|\Delta|^{2}}}-\int\frac{\mathrm{d}\epsilon}{2}\frac{\mathrm{d}\phi}{2\pi}\;\frac{n_{F}\left[{\left(\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F}\right)}/{k_{B}T}\right]}{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}.$ (9)
The first two terms in Eq. 9 give $\log(\Delta_{0}/\bar{\beta}|\Delta|)$. The integral over the Fermi-distribution on the right-hand side of Eq. 9 may be rewritten as a series expansion
$\int\frac{\mathrm{d}\phi}{2\pi}n_{F}\left[{\left(\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F}\right)}/{k_{B}T}\right]$ (10)
$=$ $\int\frac{\mathrm{d}\phi}{2\pi}\sum_{n=1}^{\infty}(-1)^{n-1}\exp\left[-n\frac{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F}}{k_{B}T}\right]$
$=$ $\sum_{n=1}^{\infty}(-1)^{n-1}\exp\left[-n\frac{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}{k_{B}T}\right]I_{0}\left(n\frac{\alpha p_{F}\bar{H}}{k_{B}T}\right).$
For $\alpha p_{F}\gg k_{B}T$ the modified Bessel function becomes $I_{0}\left({n\alpha p_{F}\bar{H}}/{k_{B}T}\right)\simeq\sqrt{{k_{B}T}/{2\pi n\alpha p_{F}\bar{H}}}\exp\left({n\alpha p_{F}\bar{H}}/{k_{B}T}\right).$ Similarly, for $|\Delta|\gg k_{B}T$ we may approximate the exponential in Eq. 10 by ${\exp\left(-n{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}/{k_{B}T}\right)}/{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}}\simeq{\exp\left(-n{(\bar{\beta}|\Delta|+\epsilon^{2}/2\bar{\beta}|\Delta|)/k_{B}T}\right)}/{\bar{\beta}|\Delta|}.$ Using these approximations, the integral over energy in Eq. 8 reads
$\int\frac{\mathrm{d}\epsilon}{2}\frac{\mathrm{d}\phi}{2\pi}\;\frac{n_{F}\left[{\left(\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}}+\alpha\nu(\mathbf{\bar{H}}\times\hat{z}).\mathbf{p}_{F}\right)}/{k_{B}T}\right]}{\sqrt{\epsilon^{2}+\bar{\beta}^{2}|\Delta|^{2}},}$ (11)
$\simeq\sum_{n=1}^{\infty}\int\frac{\mathrm{d}\epsilon}{2}\;(-1)^{n-1}\frac{\exp\left[-\frac{n\epsilon^{2}}{2\bar{\beta}|\Delta|k_{B}T}\right]}{\bar{\beta}|\Delta|}\sqrt{\frac{k_{B}T}{2\pi n|\alpha|}}\exp\left[\frac{n(\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|)}{k_{B}T}\right]$
$=\sum_{n=1}^{\infty}\frac{(-1)^{n-1}}{2n}\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\exp\left[\frac{n(\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|)}{k_{B}T}\right]$
$=\frac{1}{2}\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\log\left(1+\exp\left[\frac{\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|}{k_{B}T}\right]\right).$
Moreover, the sums over Matsubara frequencies on the right-hand hand side of Eq. 8 are known to give polygamma functions. To obtain the final expression for the self-consistency equation, we reinterpret the BCS energy cutoff $\omega_{D}$ in terms of $T_{c}$ and $\Delta_{0}$. To achieve this, we use that $1/\mathcal{N}\lambda=\log(2\omega_{D}/|\Delta_{0}|)=\sum_{n=0}^{\hbar\omega_{D}/2\pi k_{B}T_{c}}{(n+1/2)^{-1}}$ at zero field. Hence, the self-consistency equation at low temperatures $\Delta\gg k_{B}T$ simplifies to
$\log\left(\frac{|\Delta_{0}|}{\bar{\beta}|\Delta|}\right)=$ $\sum_{\nu}\frac{1}{2}\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\log\left(1+\exp\left[\frac{\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|}{k_{B}T}\right]\right)+\frac{H^{2}}{\beta^{2}}\Bigg[\log\left(\frac{T}{T_{c}}\right)-\psi\left(\frac{1}{2}\right)+\log\left(\frac{\sqrt{H^{2}+\beta^{2}}}{2\pi k_{B}T}\right)\Bigg].$ (12)
[…]
$\log\left(\frac{|\Delta_{0}|}{\bar{\beta}|\Delta|}\right)=$ $\frac{k_{B}T}{\sqrt{\bar{\beta}|\Delta|\alpha p_{F}\bar{H}}}\log\left(1+\exp\left[\frac{\alpha p_{F}\bar{H}-\bar{\beta}|\Delta|}{k_{B}T}\right]\right)+\frac{H^{2}}{\beta^{2}}\log\left(\frac{2\sqrt{H^{2}+\beta^{2}}}{\Delta_{0}}\right).$ (13)
II High temperature Ginzburg-Landau limit
To determine the high temperature limit, $(T_{c}-T)/T\ll 1$, we expand the self-consistency equation (7) up to second order in $\bar{\beta}|\Delta|$ and $\alpha p_{F}$ and ignore the contribution from the helical state $v_{F}q\rightarrow 0$. After taking the angular integral, the self-consistency equation (7) reads
$\frac{1}{\mathcal{N}\lambda}=\pi\mathrm{i}k_{B}T\sum_{n,\nu}\;\bar{\beta}^{2}\mathrm{sign}(\omega_{n})$ $\left[\frac{1}{\mathrm{i}{\omega_{n}}}-\frac{\alpha^{2}p_{F}^{2}\bar{H}^{2}+\bar{\beta}^{2}|\Delta|^{2}}{2\mathrm{i}\omega_{n}^{3}}\right]$ (14)
$+{\bar{H}^{2}\mathrm{sign}(\omega_{n})}$ $\left[\frac{1}{\mathrm{i}\omega_{n}-\nu\sqrt{H^{2}+\beta^{2}}}\right].$
The sums over Matsubara frequencies are known to give polygamma functions and the above self-consistency equation becomes
$\log\left(\frac{T}{T_{c}}\right)=\bar{H}^{2}$ $\left[\psi\left(\frac{1}{2}\right)-\mathrm{Re}\psi\left(\frac{1}{2}-\mathrm{i}\frac{\sqrt{H^{2}+\beta^{2}}}{2\pi k_{B}T}\right)\right]-\bar{\beta}^{2}\left[\frac{7\zeta(3)}{2(2\pi k_{B}T)^{2}}\left(\alpha^{2}p_{F}^{2}\bar{H}^{2}+\bar{\beta}^{2}|\Delta|^{2}\right)\right].$ (15)
Since the Ising SOC is typically much larger than the critical temperature $\sqrt{\beta^{2}+H^{2}}\gg k_{B}T$ the digamma function is well approximated by a logarithm $\mathrm{Re}\psi\left({1}/{2}-\mathrm{i}{\sqrt{H^{2}+\beta^{2}}/}{2\pi k_{B}T}\right)\rightarrow\log\left({\sqrt{H^{2}+\beta^{2}}}/{2\pi k_{B}T}\right).$ Thus, the above is well approximated by
$\log\left(\frac{T}{T_{c}}\right)=\bar{H}^{2}\left[\psi\left(\frac{1}{2}\right)-\log\left(\frac{\sqrt{H^{2}+\beta^{2}}}{2\pi k_{B}T}\right)\right]-\bar{\beta}^{2}$ $\left[\frac{7\zeta(3)}{2(2\pi k_{B}T)^{2}}\left(\alpha^{2}p_{F}^{2}\bar{H}^{2}+\bar{\beta}^{2}|\Delta|^{2}\right)\right].$ (16)
The above can be written more compactly as
$\frac{7\zeta(3)}{2(2\pi k_{B}T)^{2}}\left(\bar{\beta}^{2}|\Delta|^{2}+\alpha^{2}p_{F}^{2}\bar{H}^{2}\right)=\log\left(\frac{T_{c}}{T}\right)-\frac{H^{2}}{\beta^{2}}\log\left(\frac{2\sqrt{H^{2}+\beta^{2}}}{\Delta_{0}}\right).$ (17)
Hence, we see the SC OP does not collapse at $\beta|\Delta|=\alpha p_{F}H$ for temperatures close to $T_{c}$.

### v2

@@ Abstract
Introduction.— Ising superconductors (SC) are doped two-dimensional transition metal dichalcogenides (TMDs), which are characterized by strong resilience against in-plane magnetic fields beyond the Pauli-paramagnetic limit. As a platform, TMDs are under active interest due to their direct band gap atomic scale thickness and strong spin-orbit coupling (SOC) [19]. Like graphene, TMDs have a honeycomb lattice and hence two valleys. Unlike graphene, however, TMDs exhibit strong SOC because of their heavy atoms and in-plane inversion symmetry breaking. These features make them particularly interesting for potential applications in optoelectronics, valleytronics, and spintronics [27, 18]. When sufficiently doped, TMDs can become superconducting (SC) [14, 4, 8, 24, 10, 21, 22, 21, 20, 3]. Due to their large SOC, Cooper pairs are Ising coupled, making them resistant to large external magnetic fields [14, 4, 8, 24, 10, 21, 22, 21, 20, 3].
@@
Besides the Ising SOC, one expects a Rashba SOC to originate from gating or interfacial effects in experiment—due to the heavy atoms of the Ising SC. The effect of Rashba SOC can lead to new physics. For instance, previously, it has been shown that the presence of Rashba SOC can lead to topological superconducting phases [29, 2, 23]. Experimentally, the strength of Rashba SOC has been claimed to be of the order of the SC order parameter [16, 17], making Rashba SOC possibly important for understanding the phase diagram of Ising SCs.
@@
In this Letter, we show that Rashba SOC sets a critical magnetic field—in the $s$-wave two valley model of Ising SCs—beyond which the superconducting order parameter (SC OP) collapses. This effect occurs at low temperatures and disappears close to the critical temperature $T_{c}$. While the effect of Rashba SOC is usually subtle, e.g., in the superconducting diode effect [28, 9, 1]. This breakdown of the superconducting gap is a rather strong effect. Our results show that low and high temperature physic of SCs can behave both quantitatively and qualitatively different. Furthermore, our results may lead to new ways to determine the strength of the Rashba SOC in Ising SCs, which is currently a challenging task.
@@
The model.— The low energy physics of doped transition metal dichalcogenides is described by two valleys around the $\mathrm{K}$ and $\mathrm{K}^{\prime}=-\mathrm{K}$ points and a gamma pocket $\Gamma$. In the proceeding, we focus on the physics of the $\pm\mathrm{K}$ points and ignore the physics of the Gamma pocket. The general Hamiltonian near the $\pm\mathrm{K}$ [27, 13, 21, 10] points in the $(a_{\mathbf{p},\uparrow},a_{\mathbf{p},\downarrow})$ basis reads
@@
The off-diagonal term $\mathrm{i}\sigma_{y}\otimes\delta\check{H}$ in the inverse Green’s function above, can be treated perturbatively since $|\Delta|/\sqrt{H^{2}+\beta^{2}}$ is found to be at most of the order $\mathcal{O}(10^{-1})$ in Ising SCs [4, 7, 8, 26, 14, 17, 16]. Hence, to good accuracy we may expand the Gorkov Green’s function up to first order. The lowest order Gorkov Green’s function has a BCS form and reads
@@
$\frac{f_{\mathrm{i}\omega_{n},+,-}^{0}(\eta,\phi)}{\bar{H}\Delta}=$ $\frac{2\pi\mathrm{i}\mathcal{N}\mathrm{sign}(\omega_{n})}{\mathrm{i}\omega_{n}-\sqrt{H^{2}+\beta^{2}}},$
$\frac{f_{\mathrm{i}\omega_{n},-,+}^{0}(\eta,\phi)}{-\bar{H}\Delta}=$ $\frac{2\pi\mathrm{i}\mathcal{N}\mathrm{sign}(\omega_{n})}{\mathrm{i}\omega_{n}+\sqrt{H^{2}+\beta^{2}}}.$
@@
In the following, we discuss the effect of Rashba SOC on the magnetic field dependence of the SC OP at low temperatures. In principle, the effect of Rashba on Ising SCs is two-fold. One effect is that Rashba SOC coupling alters the thermodynamic ground state to a helical state $\Delta\rightarrow\Delta_{0}e^{\mathrm{i}qx}$ [11, 28]. This effect is however negligible with respect to the magnitude of the SC OP since $v_{F}q$ is at most of the order $\mathcal{O}(\alpha p_{F}/\beta)$ and the correction of the helical phase to $|\Delta|/\Delta_{0}$ is at most of the order $\mathcal{O}(\alpha^{2}p_{F}^{2}/\beta^{2})<\mathcal{O}(10^{-2})$. However, the main effect of Rashba SOC to the magnitude of the SC OP is that the SC OP collapses beyond a critical magnetic field set by Rashba SOC, as we will show next.
@@
In the following, we focus on low temperatures $k_{B}T\ll\alpha p_{F}\lesssim\Delta_{0}\ll\sqrt{H^{2}+\beta^{2}}$. Translated to the critical temperature, this means $T/T_{c}\lesssim 10^{-2}$. Since $T_{c}\sim\mathcal{O}(10\mathrm{K})$ in Ising SCs, this corresponds to temperatures of the order $T\lesssim\mathcal{O}(100\mathrm{mK})$ [14, 8, 15, 26]. Such temperatures are typically accessible in experiment [14]. Furthermore, we consider the energy scale of Rashba SOC to be at most of the order of the energy scale of the SC OP at zero temperature. As detailed in the supplementary material, we perform the summation over Matsubara frequencies in the first term of Eq. 3, which is known to give the digamma function [12] and approximate the second term for low temperatures. As a result, the self-consistency equation (3) for low temperatures reads
@@
In Refs. [16, 17] values of $\alpha p_{F}\sim\Delta_{0}$ in $\mathrm{MoS_{2}}$ and $\mathrm{WS_{2}}$ were reported. This should be sufficient for the collapse of the SC OP to happen at low temperatures.
@@
Hence, the effective gap that is measured via the density of states is $\bar{\beta}|\Delta|$ rather than $|\Delta|$ and Rashba SOC makes the density of states cone-shaped for energies $\epsilon$ satisfying $|\epsilon-\bar{\beta}|\Delta||<\alpha p_{F}H$, see Fig. 4. Moreover, inclusion of the second order correction in Gorkov Green’s function (2) gives what is called a mirage gap at high energies $\epsilon\sim\sqrt{\beta^{2}+H^{2}}$ in the density of states [25].
@@
Discussion.— In this Letter, we showed that the presence of Rashba SOC in Ising superconductors can strongly affect the magnetic field dependence of the superconducting order parameter at low temperatures. Our main finding is that the SC OP collapses beyond a critical magnetic field satisfying $\alpha p_{F}H=\beta|\Delta|$. This collapse disappears at high temperatures—close to $T_{c}$—and demonstrates how the low- and high temperature limits of Ising SCs can be quantitatively and qualitatively different. We find Rashba SOC to be significant to the superconducting phase diagram if $\alpha p_{F}/\Delta_{0}\gtrsim\mathcal{O}(10^{-1})$. Due to the heavy atom nature of TMDs, such a Rashba SOC could be induced from gating or interfacial effects in experiments. Experimentally, Rashba SOCs of magnitude of the SC OP have already been measured in thin-film Ising SCs [16, 17]. Our findings may lead to new ways to determine the size of Rashba SOC in Ising SCs, which is currently a challenging task.
The collapse of the SC OP at low temperatures should be experimentally verifiable using tunneling spectroscopy [14, 5, 3]. While different tunneling devices have different sensitivities towards either the $\Gamma$ pocket or the $\pm\mathrm{K}$ valleys, the predicted effect should be within experimental reach using tunneling devices that are mostly sensitive to the density of states in the $\pm\mathrm{K}$ valleys [14].
@@
In future work, one could extend this model to include the $\Gamma$ pocket [6]. Additionally, inclusion of impurities and scattering between the $\pm\mathrm{K}$ valleys and $\Gamma$ pocket would be of interest.